OpenStudy (anonymous):

MEDAL FAN AND TESTIMONIAL:on a camping trip you bring 12 items for 4 dinners for each dinner you use 3 items in how many ways can you choose the 3 items for the first dinner? for the second? for the third? for the fourth?

OpenStudy (anonymous):

its on permutations and combinations algebra 2

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

@mathstudent55 ???

OpenStudy (anonymous):

It is combinations. And the result is same for the all four dinners.

OpenStudy (anonymous):

how do i do it?

OpenStudy (anonymous):

I can do the problems that are already set up but not word problems where i have to set it up

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

You have four dinner. And for all of them, you need to make a decision to pick 3 items among 12 items. So 3 combinations among 12 item is: \[\left(\begin{matrix}12 \\ 3\end{matrix}\right) = \frac{ 12 \times 11 \times 10 }{ 9! \times 3! }\] I had a mistake though, in my first statement, where I said that they were equal. It is not. If you pick 3 items among 12, you will have 9 items left, since you took 3 of them.

OpenStudy (anonymous):

eek i got a weird answer

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

\[11/18144\]

OpenStudy (anonymous):

that was evaluating the right side

OpenStudy (anonymous):

ohh sorry, I forgot to add factorial sign to the 12. It is like this: \[\frac{ 12! }{ 9! \times 3! }\] it is originally this way.

OpenStudy (anonymous):

The general formula is: \[\left(\begin{matrix}c \\ n\end{matrix}\right) = \frac{ c! }{ n! \times (c - n)! }\]

OpenStudy (anonymous):

i got 220

OpenStudy (anonymous):

is that for the first dinner

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok how do i find the others

OpenStudy (anonymous):

For the second dinner you will calculate this: \[\left(\begin{matrix}9 \\ 3\end{matrix}\right)\] Since we took the first three items, 9 left in the item list.

OpenStudy (anonymous):

not sure how to solve that?

OpenStudy (anonymous):

its 9/3??

OpenStudy (anonymous):

Check my answer, I generalized the formula for you

OpenStudy (anonymous):

\[\left(\begin{matrix}c \\ n\end{matrix}\right) = \frac{c!}{n! \times (c-n)!}\]

OpenStudy (anonymous):

i got 84

OpenStudy (anonymous):

yeah, probably, I didn't calculated

OpenStudy (anonymous):

ok 3rd?

OpenStudy (anonymous):

How many items do we left? We used 6 of them so far right?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

yes, you will chose 3 from 6. Use the formula that I gave you. Did you understand how to use the formula?

OpenStudy (anonymous):

mm let me see ill try

OpenStudy (anonymous):

i got 20

OpenStudy (anonymous):

yes. I think

OpenStudy (anonymous):

ok for the fourth then all of the variables in the formula you gave are 3 right

OpenStudy (anonymous):

yes, you have learned ;)

OpenStudy (anonymous):

i think so but the problem is that I'm getting 1/6?

OpenStudy (anonymous):

you should get 1. Check if you didn't make any calculation mistakes.

OpenStudy (anonymous):

or i guess i would just say one because there isn't anything left

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

oh haha okay thank you so much (:

OpenStudy (anonymous):

Anytime ;)